OpenStudy (anonymous):

Write the equation (in vertex form) for the parabola with the given vertex and passing through the given point. Vertex (2,0) point (1,4)

OpenStudy (anonymous):

a parabola will have the same form as a the standard form used in the last question, like this: Ax^2 + Bx + C = 0

OpenStudy (anonymous):

alrighty but which numbers do you use for what?

OpenStudy (anonymous):

one second, i'm finding it so i teach you right..

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

ok ok, for get what I said before.. so it's like this: the vertex has the terms (h, k) so we have (2,0) so that means h = 2 k= 0 make sense?

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

now here is the form for the parabola: y = a(x-h)^2 + k we know what h and k are, and the point given to us of (1,4) we can use for x and y, so we will need to plug those numbers in and solve for a

OpenStudy (anonymous):

soo hmm (1-2)^2+0??

OpenStudy (anonymous):

dont forget the part "y = a(" and plug in y, becuase we know what that is

OpenStudy (anonymous):

never mind it says just write the equation..

OpenStudy (anonymous):

right, but we arn't done, we have to find what a is equal to to wrtie the equation

OpenStudy (anonymous):

o alright, soo hmm 4=a(1-2)^2+0

OpenStudy (anonymous):

good good, now solve for a

OpenStudy (anonymous):

is it 4?.....4=4(1-2)^2+0

OpenStudy (anonymous):

i'll walk you through it-

OpenStudy (anonymous):

4=a(1-2)^2+0 4=a(-1)^2 4=a(1) a=4, then we put it into the proper form, plug in h, k, a, but not x and y into this equation: y = a(x-h)^2 + k

OpenStudy (anonymous):

umm 4(x-2)^2+0

OpenStudy (anonymous):

very good, just don't forget the y=, so like this: y=4(x-2)^2 good job

OpenStudy (anonymous):

so what the equation means is that we can plug in any x value we want, and it will spit out the value of Y for that parabola good job i have to go for the night

OpenStudy (anonymous):

alright thanks alot